OpenStudy (anonymous):

which of the following is equivalent to the equations 11-3x^2+7x-4=8x+1 A.3x^2+15x-6=0 B.3x^2+15x+16=0 C.3x^2+x-6=0 D.-3x^2+x+6=0

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

\(\Large 11-3x^2+7x-4=8x+1\\\Large \Rightarrow 8x+1-[-3x^2+7x+(11-4)]=0\\\Large \Rightarrow 8x+1+3x^2-7x-7=0\\\Large \Rightarrow 3x^2+x-6=0\)